# A Banach *-algebra associated with the action of a group on a topological space

### v1

@@ 3. The ideal intersection property and consequences
We say that $\Sigma$ is residually topologically free if for any $G$-invariant closed subset $D$ of $X$, the restriction of $\Sigma$ to $D$ is topologically free. For brevity, we will sometimes write r.t.f. in place of residually topologically free.
@@
For abelian groups, the notions of free actions and of r.t.f. actions are equivalent. It is evident that if the action is free then it is also r.t.f. For the other direction, suppose that there is some $x\in X$ such that $G_{x}\neq\{e\}$. Take $D$ as the closure of the orbit of $x$. As $G$ is abelian, all points of the orbit of $x$ have the same stabilizer subgroup $G_{x}$. Then, given $t\in G_{x}$, for every $s\in G$ we have that $\sigma_{t}(\sigma_{s}(x))=\sigma_{s}(\sigma_{t}(x))=\sigma_{s}(x).$ This means that $t$ fixes the whole orbit of $x$, so it also fixes $D$, so the restriction of $\Sigma$ to $D$ is not topologically free.
@@
Let us fix some notation. Suppose that $D$ is a closed $G$-invariant subset of $X$. Then, we can restrict to $D$ the action of $G$ and obtain a new dynamical system, that we will denote $\Sigma_{D}$. As $\Sigma_{D}$ is a dynamical system, we can consider the algebra $\ell^{1}(\Sigma_{D}),$ and the map $R_{D}:\ell^{1}(\Sigma)\to\ell^{1}(\Sigma_{D})$ given by
$R_{D}\left(\sum_{s\in G}f_{s}\delta^{s}\right)=\sum_{s\in G}f_{s}|_{D}\delta^{s}_{D}\text{ for every }\sum_{s\in G}f_{s}\delta^{s}\in\ell^{1}(\Sigma),$
where $\delta_{D}^{s}$ denotes the canonical unitary element of $\ell^{1}(\Sigma_{D})$ associated to $s\in G$. It is clear that $R_{D}$ is surjective and norm-one *-homomorphism, and $\ker(R_{D})=\{\sum_{s\in G}f_{s}\delta^{s}:f_{s}|_{D}=0\text{ for every }s\in G\}$.
@@
Let $I$ be a closed ideal of $\ell^{1}(\Sigma)$, and let $D$ be a closed $G$-invariant subset of $X$ such that $I\cap C(X)=\{f\in C(X):f|_{D}=0\}$ (as in Proposition 3.12). Then $R_{D}(I)$ has zero intersection with $C(D)$.
@@
Let $g$ be in $R_{D}(I)\cap C(D)$, so there is some $\sum_{s\in G}f_{s}\delta^{s}\in I$ that gets mapped to $g$ through $R_{D}.$ We have then that
$\sum_{s\in G}f_{s}|_{D}\delta^{s}_{D}=g,$
so it follows that $f_{s}|_{D}=0$ for every $s\in G\setminus\{e\}$ and $f_{e}|_{D}=g$. We have then that $\sum_{s\neq e}f_{s}\delta^{s}\in\ker(R_{D})$. As $\{f\in C(X):f|_{D}=0\}=I\cap C(X)\subset I$, it’s clear that $\ker(R_{D})\subset I$, so it follows that
$f_{e}=\sum_{s\in G}f_{s}\delta^{s}-\sum_{s\neq e}f_{S}\delta^{s}\in I.$
Then $f_{e}\in I\cap C(X)$, so $f_{e}|_{D}=0.$ As $g=f_{e}|_{D}$, we conclude that $R_{D}(I)\cap C(D)=\{0\}$. ∎
@@
Given a bounded and surjective linear operator $T:A\to B$ between Banach spaces, and given $I$ a nonempty subset of $A$, we have that $T(I)$ is closed in $B$ if and only if $I+\ker(T)$ is closed in $A$. We will use this in the following theorem.
In the next proof, given $D$ a closed $G$-invariant subset of $X$, we denote $r_{D}:C(X)\to C(D)$ the map given by $r_{D}(f)=f|_{D}$ for every $f\in C(X)$. We also denote $E_{D}:\ell^{1}(\Sigma_{D})\to C(D)$ the map given by $E_{D}(\sum_{s\in G}f_{s}\delta^{s}_{D})=f_{e}$ for every $\sum_{s\in G}f_{s}\delta^{s}_{D}\in\ell^{1}(\Sigma_{D})$. It is clear that $E_{D}\circ R_{D}=r_{D}\circ E$.
@@
    For every pair of ideals $I$ and $J$ of $\ell^{1}(\Sigma)$, $I\cap C(X)=J\cap C(X)$ implies $I=J$.
@@
    For every closed ideal $I$ of $\ell^{1}(\Sigma)$ we have that $E(I)$ is contained on $I$.
@@
    $\Sigma$ is residually topologically free.
@@
Let’s first prove that (1) implies (2). Let $I$ be a closed ideal of $\ell^{1}(\Sigma)$. Using Proposition 3.12 we have that there is some closed and $G$-invariant subset $D$ of $X$ such that $I\cap C(X)$ consists of the countinuous functions that vanish on $D$. We define the closed ideal
$J=\left\{\sum_{s\in G}f_{s}\delta^{s}:f_{s}|_{D}=0\text{ for every }s\in G\right\}.$
It is clear that $J\cap C(X)=I\cap C(X)$, so using (1) we have that $J=I$. Now it is clear that $E(I)=I\cap C(X)\subset I$.
We will prove now that (2) implies (3). Let $D$ be a closed and $G$-invariant subset of $X$. In order to prove that $\Sigma_{D}$ is topologically free, we will use Theorem 3.10. So, let $J$ be a closed ideal of $\ell^{1}(\Sigma_{D})$ such that $J\cap C(D)=\{0\}$, we have to prove that $J=\{0\}$. We define $I=R_{D}^{-1}(J)$, which is a closed ideal of $\ell^{1}(\Sigma)$. By (2) we have that $E(I)\subset I$, so in particular $E(I)\subset I\cap C(X)$. Now, we have that
@@
As $E_{D}(J)=\{0\}$ we have that $J=\{0\}$.
That (3) implies (1) is the only implication left to prove. Suppose that $I$ and $J$ are two closed ideals of $\ell^{1}(\Sigma)$ such that $I\cap C(X)=J\cap C(X)$. Using Proposition 3.12 we get a closed and $G$-invariant subset $D$ of $X$ such that $I\cap C(X)=J\cap C(X)=\{f\in C(X):f|_{D}=0\}.$ We have that $\ker(R_{D})=\{\sum_{s\in G}f_{s}\delta^{s}:f_{s}|_{D}=0\text{ for every }s\in G\}$ is included in $I$ and $J$, and as $R_{D}$ is contractive and surjective and $I,J$ are closed, we have that, by Remark 3.21, $R_{D}(I)$ and $R_{D}(J)$ are closed ideals of $\ell^{1}(\Sigma_{D})$. Now, by Lemma 3.20 we have that $R_{D}(I)\cap C(D)=R_{D}(I)\cap C(D)=\{0\}$. By (3) we have that $\Sigma_{D}$ is topologically free, so by Theorem 3.10 we have $R_{D}(I)=R_{D}(J)=0$. This means that $I$ and $J$ are included in $\ker(R_{D})$. As we have proved both inclusions hold, we have that $I=\ker(R_{D})=J$. ∎
@@
The theorem above bears a resemblance to some of the work done in [dJT13, Section 3]. There, for $\mathbb{Z}$-actions, two notions of noncommutative spectral synthesis are studied for $\ell^{1}(\Sigma)$. It is proved there that (1) and (2) are equivalent to the action being free, which makes sense given that $\mathbb{Z}$ is abelian and Remark 3.18.
@@ 4. Detecting freeness through self-adjoint ideals
We mentioned in the introduction the result [dJST12, Theorem 4.4]. This theorem states that, for $\mathbb{Z}$-actions, the action is free if and only if every closed ideal of $\ell^{1}(\Sigma)$ is self-adjoint. This theorem is interesting in the sense that freeness of the action can be detected by $\ell^{1}(\Sigma)$ through the property of every closed ideal being automatically self-adjoint. This is not something that can be detected by its $C^{*}$-envelope, as in a $C^{*}$-algebra all closed ideals are self-adjoint. We would like to generalize this result to this setting, but some problems arise when trying to do this. First of all, the following example shows that [dJST12, Theorem 4.4] is not true in general.
@@
We have an action $\varphi:\mathbb{Z}_{2}\to\text{Aut}(\mathbb{Z})$ of $\mathbb{Z}_{2}$ on $\mathbb{Z}$, given by $\varphi(0)=\text{Id}$ and $\varphi(1)=-\text{Id}$, so we can consider the semidirect product $G=\mathbb{Z}\rtimes_{\varphi}\mathbb{Z}_{2}$. Let $\theta$ be an irrational number. Now, let $X$ be the circle $S^{1}=\mathbb{R}/\mathbb{Z}$, and let $G$ act on $X$ the following way: $\sigma_{(n,0)}(\beta)=\beta+n\theta$ (modulo $\mathbb{Z}$) and $\sigma_{(0,1)}(\beta)=-\beta$ for every $\beta\in\mathbb{R}/\mathbb{Z}$. It is direct to check that this defines an action on $X$, given by $\sigma_{(n,m)}(\beta)=(-1)^{m}\beta+n\theta$. If we identify $\mathbb{R}/\mathbb{Z}$ with the unit circle on the complex plane, what we have is that $(1,0)$ acts as an irrational rotation of angle $\theta$ and that $(0,1)$ acts as a symmetry with respect to the real axis. As $\theta$ is irrational we have that $\Sigma$ is minimal. It is easy to check that $\text{Per}(\Sigma)=\left\{\beta=\frac{n\theta+m}{2}\in\mathbb{R}/\mathbb{Z}:n,m\in\mathbb{Z}\right\}$, so the action is topologically free but not free. But as the action is topologically free and minimal, by Theorem 3.13 we have that $\ell^{1}(\Sigma)$ has no proper closed ideals, so trivially every closed ideal of $\ell^{1}(\Sigma)$ is self-adjoint.
We remark that the proof of one of the implications of [dJST12, Theorem 4.4] generalizes well to this setting: if $\Sigma$ is free, then every closed ideal of $\ell^{1}(\Sigma)$ is self-adjoint. It is interesting that, in that proof, what is actually used is that if the action is free then it is also r.t.f, and that implies that every closed ideal is self-adjoint. This is not a big deal if we are working with $\mathbb{Z}$-actions because of Remark 3.18, but if we are working with a nonabelian group, it tells us that that being r.t.f, which is weaker than the action being free, is enough to have that every closed ideal is self-adjoint. Regardless, we are able to obtain this implication easily as a simple corollary of Theorem 3.22.
@@
If $\Sigma$ is residually topologically free, then every closed ideal of $\ell^{1}(\Sigma)$ is self-adjoint.
@@
Let $I$ be a closed ideal. $C(X)$ is a $C^{*}$-algebra, so every closed ideal is self-adjoint, so it must be that $I\cap C(X)=I^{*}\cap C(X)$. As $\Sigma$ is r.t.f, by Theorem 3.22 we obtain that $I=I^{*}$. ∎
As we said before, freeness of the action is stronger than it being r.t.f, so the above theorem also works if we change one condition by the other. We will now try to generalize the reciprocal of the theorem. We cannot do the same that was done in [dJST12, Theorem 4.4] because there it is used that if the action of $\mathbb{Z}$ is not free then there is a finite orbit, which is not necessarily true for arbitrary discrete and countable groups.
We denote by $\ker(\alpha)$ the set of elements of $G$ that act trivially on $X$. In what follows, given a discrete group $H$, we denote by $\ell^{1}(H)$ the Banach *-algebra of integrable functions with respect to the counting measure, with the usual convolution and involution, and we denote its norm by $\lVert\>\cdot\>\rVert_{1}$. Given $t\in H$ we write $\mathbbm{1}_{t}\in\ell^{1}(H)$ the characteristic function of $\{t\}$, and we recall that $(\mathbbm{1}_{t}*\xi)(h)=\xi(t^{-1}h)$ for every $\xi\in\ell^{1}(H)$ and $h\in H$, where * denotes the usual convolution in $\ell^{1}(H)$.
@@
Suppose that $G$ is abelian. Let $H$ be a subgroup of $G$ contained on $\ker(\alpha)$ and such that $\ell^{1}(H)$ has a closed and non-self-adjoint ideal. Then $\ell^{1}(\Sigma)$ contains a closed and non-self-adjoint ideal.
@@
Let $J$ be a closed and non-self-adjoint ideal of $\ell^{1}(H)$. For every $x\in X,s\in G$ and $f\in\ell^{1}(\Sigma)$ we define $\phi_{x,s}(f)$ as the map from $H$ to $\mathbb{C}$ given by $\phi_{x,s}(f)(h)=f(sh)(x)$ for every $h\in H$. We have that
$\begin{split}\lVert\phi_{x,s}(f)\rVert_{1}=\sum_{t\in H}|\phi_{x,s}(f)(t)|=\sum_{t\in H}|f(st)(x)|\leq\sum_{t\in H}\lVert f(st)\rVert_{\infty}\leq\sum_{s\in G}\lVert f(s)\rVert_{\infty}=\lVert f\rVert<\infty,\end{split}$
so $\phi_{x,s}(f)$ belongs to $\ell^{1}(H)$ for every $x\in X,s\in G$ and $f\in\ell^{1}(\Sigma)$. We can then think of $\phi_{x,s}$ as a contracting linear map from $\ell^{1}(\Sigma)$ to $\ell^{1}(H)$. We define
$\widetilde{J}=\left\{f\in\ell^{1}(\Sigma):\phi_{x,s}(f)\in J\text{ for every }x\in X\text{ and }s\in G\right\}.$
We will prove that $\widetilde{J}$ is closed and non-self-adjoint ideal of $\ell^{1}(\Sigma)$.
@@
$\widetilde{J}=\bigcap_{\begin{subarray}{c}x\in X\\ s\in G\end{subarray}}\phi_{x,s}^{-1}(J),$
and as for every $x\in X$ and $s\in G$ we have that $\phi_{x,s}$ is contractive and $J$ is closed, we have that $\widetilde{J}$ is also closed. We have to check that it is an ideal. We write $G/H=\{s_{\beta}H:\beta\in I_{H}\}$ where $\{s_{\beta}:\beta\in I_{H}\}$ is a set of representatives of the quotient, so we have that $G=\bigsqcup_{\beta\in I_{H}}s_{\beta}H$. Let $f\in\ell^{1}(\Sigma)$ and $g\in\widetilde{J}$, and fix $x\in X$ and $s\in G$. Let $h\in H$, we have that

### v2

@@ 3. The ideal intersection property and consequences
We say that the action $G\curvearrowright X$ is residually topologically free if for any $G$-invariant closed subset $D$ of $X$, the restriction of the action of $G$ to $D$ is topologically free. For brevity, we sometimes write r.t.f. in place of residually topologically free.
@@
For abelian groups, the notions of free actions and of r.t.f. actions are equivalent. It is evident that if the action is free then it is also r.t.f. For the other direction, suppose that there is some $x\in X$ such that $G_{x}\neq\{e\}$. Take $D$ as the closure of the orbit of $x$. As $G$ is abelian, all points of the orbit of $x$ have the same stabilizer subgroup $G_{x}$. Then, given $t\in G_{x}$, for every $s\in G$ we have that $\sigma_{t}(\sigma_{s}(x))=\sigma_{s}(\sigma_{t}(x))=\sigma_{s}(x).$ This means that $t$ fixes the whole orbit of $x$, so it also fixes $D$, so the restriction of the action to $D$ is not topologically free.
@@
Let us fix some notation. Suppose that $D$ is a closed $G$-invariant subset of $X$. Then, we can restrict to $D$ the action of $G$ and obtain a new action, that we denote by $G\curvearrowright D$. As $G\curvearrowright D$ is an action, we can consider the algebra $\ell^{1}(G\curvearrowright D),$ and the map $R_{D}:\ell^{1}(G\curvearrowright X)\to\ell^{1}(G\curvearrowright D)$ given by
$R_{D}\left(\sum_{s\in G}f_{s}\delta_{s}\right)=\sum_{s\in G}f_{s}|_{D}\delta_{s}\text{ for every }\sum_{s\in G}f_{s}\delta_{s}\in\ell^{1}(G\curvearrowright X);$
here, $\delta_{s}$ also denotes the canonical unitary element of $\ell^{1}(G\curvearrowright D)$ associated to each $s\in G$. It is clear that $R_{D}$ is a surjective and norm-one *-homomorphism, and that $\ker(R_{D})=\{\sum_{s\in G}f_{s}\delta_{s}:f_{s}|_{D}=0\text{ for every }s\in G\}$.
@@
Let $I$ be a closed ideal of $\ell^{1}(G\curvearrowright X)$, and let $D$ be a closed $G$-invariant subset of $X$ such that $I\cap C(X)=\{f\in C(X):f|_{D}=0\}$ (as in Proposition 3.9). Then $R_{D}(I)$ has zero intersection with $C(D)$.
@@
Let $g\in R_{D}(I)\cap C(D)$, which means that there is some $\sum_{s\in G}f_{s}\delta_{s}\in I$ that gets mapped to $g$ through $R_{D}.$ We have then that
$\sum_{s\in G}f_{s}|_{D}\delta_{s}=g,$
so it follows that $f_{s}|_{D}=0$ for every $s\in G\setminus\{e\}$ and $f_{e}|_{D}=g$. This implies that $\sum_{s\neq e}f_{s}\delta_{s}\in\ker(R_{D})$. As $\{f\in C(X):f|_{D}=0\}$ is equal to $I\cap C(X)$, which in turn is contained in $I$, it’s clear that $\ker(R_{D})\subset I$, so it follows that
$f_{e}=\sum_{s\in G}f_{s}\delta_{s}-\sum_{s\neq e}f_{s}\delta_{s}\in I.$
Then $f_{e}\in I\cap C(X)$, so its restriction to $D$ is equal to the zero function. As $g$ is equal to $f_{e}|_{D}$, we conclude that $R_{D}(I)\cap C(D)=\{0\}$. ∎
@@
Given a bounded and surjective linear operator $T:A\to B$ between Banach spaces, and given $I$ a nonempty subset of $A$, we have that $T(I)$ is closed in $B$ if and only if $I+\ker(T)$ is closed in $A$. We use this in the following theorem.
In the next proof, given $D$ a closed $G$-invariant subset of $X$, we denote by $r_{D}:C(X)\to C(D)$ the map given by $r_{D}(f)=f|_{D}$ for every $f\in C(X)$. We also denote by $E_{D}:\ell^{1}(G\curvearrowright D)\to C(D)$ the map given by $E_{D}(\sum_{s\in G}f_{s}\delta_{s})=f_{e}$ for every $\sum_{s\in G}f_{s}\delta_{s}\in\ell^{1}(G\curvearrowright D)$. It is clear that $E_{D}\circ R_{D}=r_{D}\circ E$.
@@
    For every pair of closed ideals $I$ and $J$ of $\ell^{1}(G\curvearrowright X)$, if $I\cap C(X)$ is equal to $J\cap C(X)$ then $I=J$.
@@
    $E(I)$ is contained in $I$ for every closed ideal $I$ of $\ell^{1}(G\curvearrowright X)$.
@@
    The action $G\curvearrowright X$ is residually topologically free.
@@
Let’s first prove that (1) implies (2). Let $I$ be a closed ideal of $\ell^{1}(G\curvearrowright X)$. Using Proposition 3.9 we have that there is some closed and $G$-invariant subset $D$ of $X$ such that $I\cap C(X)$ consists of the continuous functions that vanish on $D$. We define the closed ideal
$J=\left\{\sum_{s\in G}f_{s}\delta_{s}:f_{s}|_{D}=0\text{ for every }s\in G\right\}.$
It is clear that $J\cap C(X)=I\cap C(X)$, so using (1) we have that $J=I$. Now it follows that $E(I)=I\cap C(X)\subset I$.
We prove now that (2) implies (3). Let $D$ be a closed and $G$-invariant subset of $X$, and consider the restriction of the action to $D$. In order to prove that $G\curvearrowright D$ is topologically free, we use Theorem 3.7. Let $J$ be a closed ideal of $\ell^{1}(G\curvearrowright D)$ such that $J\cap C(D)=\{0\}$, we have to prove that $J=\{0\}$. We define $I=R_{D}^{-1}(J)$, which is a closed ideal of $\ell^{1}(G\curvearrowright X)$. By (2) we have that $E(I)\subset I$, so in particular $E(I)\subset I\cap C(X)$. Now, we have that
@@
As $E_{D}(J)=\{0\}$ it follows that $J=\{0\}$.
That (3) follows from (1) is the only implication left to prove. Suppose that $I$ and $J$ are two closed ideals of $\ell^{1}(G\curvearrowright X)$ such that $I\cap C(X)=J\cap C(X)$. Using Proposition 3.9 we get a closed and $G$-invariant subset $D$ of $X$ such that both $I\cap C(X)$ and $J\cap C(X)$ are equal to $\{f\in C(X):f|_{D}=0\}.$ We have that
$\ker(R_{D})=\left\{\sum_{s\in G}f_{s}\delta_{s}:f_{s}|_{D}=0\text{ for every }s\in G\right\}$
is included in $I$ and $J$. As $R_{D}$ is contractive and surjective, and $I$ and $J$ are closed, we have by Remark 3.18 that both $R_{D}(I)$ and $R_{D}(J)$ are closed ideals of $\ell^{1}(G\curvearrowright D)$. Now, using Lemma 3.17 we get that $R_{D}(I)\cap C(D)$ and $R_{D}(I)\cap C(D)$ are both equal to $\{0\}$. By (3), we have that $G\curvearrowright D$ is topologically free, so Theorem 3.7 implies that $R_{D}(I)=R_{D}(J)=0$. This means that $I$ and $J$ are included in $\ker(R_{D})$. As we have proved both inclusions hold, we conclude that $I=\ker(R_{D})=J$. ∎
@@
The theorem above bears a resemblance to some of the work done in [dJT13, Section 3]. There, for $\mathbb{Z}$-actions, two notions of noncommutative spectral synthesis are studied for $\ell^{1}(G\curvearrowright X)$. It is proved there that (1) and (2) of the above theorem are equivalent to the action being free, which makes sense given that $\mathbb{Z}$ is abelian and Remark 3.15.
Remark 3.21.
Theorem 3.19 implies that, when the action $G\curvearrowright X$ is r.t.f., then we can completely describe the (closed) ideal structure of $\ell^{1}(G\curvearrowright X)$. Concretely, we have that every ideal closed $I$ is such that there exists a $G$-invariant closed subset $D$ of $X$ that satisfies
$I=\left\{\sum_{s\in G}f_{s}\delta_{s}:f_{s}|_{D}=0\text{ for every }s\in G\right\}.$
In this situation, we refer to $I$ as the ideal generated by $D$. It follows that, when the action is residually topologically free, there is a bijection between the closed ideals of $\ell^{1}(G\curvearrowright X)$ and the $G$-invariant closed subsets of $X$.
An ideal $I$ is prime if it is not the whole algebra and $JK\subset I$ implies $J\subset I$ or $K\subset I$, for every pair of ideals $J$ and $K$. An ideal $I$ is called semiprime if $J^{2}\subset I$ implies $J\subset I$. Equivalently, an ideal is semiprime if it is an intersection of prime ideals (see [Lam91, Theorem 10.11]). In $C^{*}$-algebras, every closed ideal is semiprime, given that every closed ideal is the intersection of primitive ideals and primitive ideals are prime. Recently, there has been an interest on the study of non-closed prime and semiprime ideals in $C^{*}$-algebras, see [GT24, GKT24, GKT25]. With the above remark at our disposal we are able to show that, if the action $G\curvearrowright X$ is r.t.f., then every closed ideal of $\ell^{1}(G\curvearrowright X)$ is semiprime. We can also characterize closed prime ideals of $\ell^{1}(G\curvearrowright X)$.
We recall that a minimal set $C$ of the action $G\curvearrowright X$ is a closed and $G$-invariant subset of $X$ such that it does not contain any proper closed and $G$-invariant subset. This is equivalent to the restriction of the action to $C$ being minimal.
Corollary 3.22.
Let $G\curvearrowright X$ be residually topologically free. Then, every closed ideal of $\ell^{1}(G\curvearrowright X)$ is semiprime. Furthermore, a closed ideal $I$ of $\ell^{1}(G\curvearrowright X)$ is prime if and only if there exists some closed and minimal $G$-invariant subset $C$ of $X$ such that
$I=\left\{\sum_{s\in G}f_{s}\delta_{s}\in\ell^{1}(G\curvearrowright X):f_{s}|_{C}=0\text{ for every }s\in G\right\}.$
Proof.
In order to check that a closed ideal $I$ is prime, it is enough to check that the condition for primeness holds with respect to closed ideals. That is, if $J$ and $K$ are closed ideals such that $JK\subset I$, then $J\subset I$ or $K\subset I$. This follows from the fact that the product ideal $\overline{J}\,\overline{K}$ is contained in $\overline{JK}$ and that, if $I$ is closed, then $\overline{I}=I$. The same argument justifies that, in order to check that a closed ideal $I$ is semiprime, it is enough to prove that if $J$ is a closed ideal such that $J^{2}$ is contained in $I$ then $J$ is included in $I$.
Let $I$ be a closed ideal and, as the action is r.t.f., let $C$ be the closed $G$-invariant subset of $X$ such that $I$ is generated by $C$, i.e.:
$I=\left\{\sum_{s\in G}f_{s}\delta_{s}:f_{s}|_{C}=0\text{ for every }s\in G\right\}.$
For every closed $G$-invariant subset $D$ of $X$, we denote by $J_{D}$ the closed ideal of $\ell^{1}(G\curvearrowright X)$ generated by $D$, which is defined analogously to $I$. It is clear that, for every pair of closed $G$-invariant sets $E$ and $D$, if $C$ is contained in $E\cup D$ then $J_{E}J_{D}$ is contained in $I$. The reciprocal also holds, that is, if $J_{E}J_{D}$ is contained on $I$, then $C$ is contained in $E\cup D$. This follows from the fact that if $J_{E}J_{D}$ is contained in $I$, then for every $f$ and $g$ belonging to $C(X)$ and vanishing on $E$ and $D$, respectively, the product $fg$ belongs to $I$, so it must vanish on $C$. This happens only if $C$ is contained in $D\cup E$. So, we have that $J_{D}J_{E}$ is contained in $I$ if and only if $C$ is contained in $D\cup E$.
Let us observe that $J_{D}$ is contained in $I$, for some closed $G$-invariant set $D$, if and only if $C$ is contained in $D$. Now, as the action is residually topologically free, every closed ideal of $\ell^{1}(G\curvearrowright X)$ is of the form $J_{D}$ for some closed $G$-invariant set $D$. This means that we can easily translate the ideal $I$ being semiprime or prime to a condition on the closed $G$-invariants sets of the action. First, $I$ is always semiprime because $J_{D}^{2}$ being contained in $I$ is equivalent to $C$ being contained in $D$, which implies that $J_{D}$ is contained in $I$. Finally, the ideal $I$ is prime if and only if for every pair $D$ and $E$ of closed $G$-invariants subsets of $X$ such that $C$ is contained in $D\cup E$ we have that $C$ is contained in $D$ or $C$ is contained in $E$. This is equivalent to $C$ being a minimal set for the action. ∎
@@ 4. Detecting freeness through self-adjoint ideals
We mentioned in the introduction the result [dJST12, Theorem 4.4]. This theorem states that, for $\mathbb{Z}$-actions, the action is free if and only if every closed ideal of $\ell^{1}(\mathbb{Z}\curvearrowright X)$ is self-adjoint. This theorem is interesting in the sense that freeness of the action can be detected by $\ell^{1}(\mathbb{Z}\curvearrowright X)$ through the property of every closed ideal being automatically self-adjoint. This is not something that can be detected by its $C^{*}$-envelope, as in a $C^{*}$-algebra all closed ideals are self-adjoint. We would like to generalize this result to this setting, but some problems arise when trying to do this. First of all, the following example shows that [dJST12, Theorem 4.4] is not true in general.
@@
We have an action $\varphi:\mathbb{Z}_{2}\to\text{Aut}(\mathbb{Z})$ of $\mathbb{Z}_{2}$ on $\mathbb{Z}$, given by $\varphi(0)=\text{Id}$ and $\varphi(1)=-\text{Id}$, so we can consider the semidirect product $G=\mathbb{Z}\rtimes_{\varphi}\mathbb{Z}_{2}$. Let $\theta$ be an irrational number. Now, let $X$ be the circle $S^{1}=\mathbb{R}/\mathbb{Z}$, and let $G$ act on $X$ the following way: $\sigma_{(n,0)}(\beta)=\beta+n\theta$ (modulo $\mathbb{Z}$) and $\sigma_{(0,1)}(\beta)=-\beta$ for every $\beta\in\mathbb{R}/\mathbb{Z}$. It is straightforward to check that this defines an action on $X$, given by
$\sigma_{(n,m)}(\beta)=(-1)^{m}\beta+n\theta.$
If we identify $\mathbb{R}/\mathbb{Z}$ with the unit circle on the complex plane, what we have is that $(1,0)$ acts as an irrational rotation of angle $\theta$ and that $(0,1)$ acts as a symmetry with respect to the real axis. As $\theta$ is irrational we have that the action $G\curvearrowright X$ is minimal. It is easy to verify that $\text{Per}(G\curvearrowright X)=\left\{\beta=\frac{n\theta+m}{2}\in\mathbb{R}/\mathbb{Z}:n,m\in\mathbb{Z}\right\}$, so the action is topologically free but not free. But as the action is topologically free and minimal, by Theorem 3.10 we have that $\ell^{1}(G\curvearrowright X)$ has no proper closed ideals so, trivially, every closed ideal of $\ell^{1}(G\curvearrowright X)$ is self-adjoint.
We remark that the proof of one of the implications of [dJST12, Theorem 4.4] generalizes well to this setting: if the action $G\curvearrowright X$ is free, then every closed ideal of $\ell^{1}(G\curvearrowright X)$ is self-adjoint. It is interesting that, in that proof, what is actually used is that if the action is free then it is also r.t.f, and that implies that every closed ideal is self-adjoint. This is not a big deal if we are working with $\mathbb{Z}$-actions because Remark 3.15 tells us that freeness and residual topological freeness are equivalent for abelian groups. But if we are working with a nonabelian group, we get that being r.t.f, which is weaker than the action being free, is enough to have that every closed ideal is self-adjoint. Regardless, we are able to prove this implication easily as a simple corollary of Theorem 3.19.
@@
If the action $G\curvearrowright X$ is residually topologically free, then every closed ideal of $\ell^{1}(G\curvearrowright X)$ is self-adjoint.
@@
Let $I$ be a closed ideal. As $C(X)$ is a $C^{*}$-algebra every closed ideal of it is self-adjoint, so it must be that $I\cap C(X)=I^{*}\cap C(X)$. As $G\curvearrowright X$ is r.t.f, by Theorem 3.19 we obtain that $I=I^{*}$. ∎
As we said before, freeness of the action is stronger than it being r.t.f, so the above theorem also works if we change one condition by the other. We now try to generalize the reciprocal of the theorem. We cannot do the same that was done in [dJST12, Theorem 4.4] because there it is used that if the action of $\mathbb{Z}$ is not free then there is a finite orbit, which is not necessarily true for arbitrary discrete and countable groups.
We denote by $\ker(\alpha)$ the set of elements of $G$ that act trivially on $X$. In what follows, given a discrete group $H$, we denote by $\ell^{1}(H)$ the Banach *-algebra of integrable functions with respect to the counting measure, with the usual convolution and involution, and we denote its norm by $\lVert\>\cdot\>\rVert_{1}$. Given $t\in H$ we denote by $\mathbbm{1}_{t}$ the characteristic function of $\{t\}$, which belongs to $\ell^{1}(H)$, and we recall that $(\mathbbm{1}_{t}*\xi)(h)=\xi(t^{-1}h)$ for every $\xi\in\ell^{1}(H)$ and every $h\in H$, where * denotes the usual convolution in $\ell^{1}(H)$.
@@
Let $G$ be abelian. Suppose that $H$ is a subgroup of $G$ contained on $\ker(\alpha)$ and such that $\ell^{1}(H)$ has a closed and non-self-adjoint ideal. Then, there is a closed and non-self-adjoint ideal in $\ell^{1}(G\curvearrowright X)$.
@@
Let $J$ be a closed and non-self-adjoint ideal of $\ell^{1}(H)$. For every $x\in X$, every $s\in G$ and every $f\in\ell^{1}(G\curvearrowright X)$, we define $\phi_{x,s}(f)$ as the map from $H$ to $\mathbb{C}$ given by $\phi_{x,s}(f)(h)=f(sh)(x)$ for every $h\in H$. We have that
$\lVert\phi_{x,s}(f)\rVert_{1}=\sum_{t\in H}|f(st)(x)|\leq\sum_{t\in H}\lVert f(st)\rVert_{\infty}\leq\sum_{s\in G}\lVert f(s)\rVert_{\infty}=\lVert f\rVert<\infty,$
so $\phi_{x,s}(f)$ belongs to $\ell^{1}(H)$ for every $x\in X$, every $s\in G$ and every $f\in\ell^{1}(G\curvearrowright X)$. Then, we can think of $\phi_{x,s}$ as a contracting linear map from $\ell^{1}(G\curvearrowright X)$ to $\ell^{1}(H)$. We define $\widetilde{J}=\left\{f\in\ell^{1}(G\curvearrowright X):\phi_{x,s}(f)\in J\text{ for every }x\in X\text{ and }s\in G\right\}.$ We will prove that $\widetilde{J}$ is a closed and non-self-adjoint ideal of $\ell^{1}(G\curvearrowright X)$.
@@
$\widetilde{J}=\bigcap_{\begin{subarray}{c}x\in X\\ s\in G\end{subarray}}\phi_{x,s}^{-1}(J)$
and, as for every $x\in X$ and every $s\in G$ we have that $\phi_{x,s}$ is contractive and $J$ is closed, we have that $\widetilde{J}$ is also closed. We have to check that it is an ideal. We write $G/H=\{s_{\beta}H:\beta\in I_{H}\}$ where $\{s_{\beta}:\beta\in I_{H}\}$ is a set of representatives of the quotient, so we have that $G=\bigsqcup_{\beta\in I_{H}}s_{\beta}H$. Let $f\in\ell^{1}(G\curvearrowright X)$ and let $g\in\widetilde{J}$, and fix $x\in X$ and $s\in G$. Let $h\in H$, we have that